OpenStudy (anonymous):

solve for x (x+1)(x+3)(x+5)(x+7) = 1920

OpenStudy (kira_yamato):

(x+1)(x+3)(x+5)(x+7) = 1920 (x^2 + 4x + 3)(x^2 + 12x + 35) = 1920 x^4 + 12x^3 + 4x^3 + 35x^2 + 3x^2 + 48x^2 + 140x + 36x + 105 = 1920 x^4 + 16x^3 + 86x^2 + 176x - 1815 = 0

OpenStudy (anonymous):

what ? so what's the answer then ... please teach me

OpenStudy (kira_yamato):

The answer is in the screenshots. x = 3 or x = -11 (if you don't consider complex no)

OpenStudy (anonymous):

x=3,−11,−4+i√39,−4−i√39

OpenStudy (anonymous):

ah, sorry forgot to saying that the x must be positive integer. hmmm, what method is it ?